OpenStudy (anonymous):

Solve for x: |3x + 2| = 2

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

You want to set it up as two different problems since it is an absolute value equation and therefore can have two different answers. Do that, and you have: 3x+2=2 or 3x+2=-2. Solve for both and get x=0 for one and x=-4/3 for the other.